Solve the simultaneous equation :f-2g+3=2f-3g+2=1
Answer by ankor@dixie-net.com(22740)

Solve the simultaneous equation :
Assume the equations are: (some commas, and/or spaces would help)
f - 2g + 3 = 2
f - 3g + 2 = 1
combine the numerical terms on the right
f - 2g = 2 - 3
f - 2g = -1
and
f - 3g = 1 - 2
f - 3g = -1
:
Using elimination
f - 2g = -1
f - 3g = -1
--------------
Subtraction eliminates f, find g
g = 0
then obviously
f = -1